post 5 - Nutritional energy

Wrong- 248 as well as 250 is wrong

± Nutritional Energy

Alexander, who weighs 189 lb , decides to climb Mt. Krumpett, which is 5560 m high. For his food supply, he decides to take nutrition bars. The label on the bars states that each 100-g bar contains 10 g of fat, 40 g of protein, and 50 g of carbohydrates.

One gram of fat contains 9 Calories, whereas each gram of protein and carbohydrates contains 4 Calories.

Nutrition facts Calories per gram
Fat 9
Protein 4
Carbohydrates 4

Part A

Alexander wants to know exactly how many bars to pack in his backpack for the journey. To provide a margin of safety, he assumes that he will need as much energy for the return trip as for the uphill climb. How many bars should Alexander pack?

Enter the exact answer to two significant figures. Do not round to an integer.
Answer #1

According to the problem Alexander's mass= 189 lb

First convert into kg as follows:

(189 lb)/(2.20462 lb/kg) = 85.7 kg (3 significant figures).

Work done = force*distance.
Force = mg =85.7 *9.81 = 840.72 N
work done = 840.72 *5560 =4674.39 kJ

Double by 2 for the return trip= 9348.78 KJ


1 calorie = 4.18 J
The energy content of each bar is:
(10 g fat)(9 Cal/g fat) = 90 Cal
(40 g protein)(4 Cal/g protein) = 160 Cal
(50 g carb)(4 Cal/g carb) = 200 Cal
Total: 450 Cal.

1 Cal = 4 186 J, so each bar gives (450 Cal)(4 186 J/Cal)

= 1883.7 K J.

Divide this to calculate ethe number of bars:
(9348.78 KJ) / (1883.7 K J. /bar)

= 4.96 bars. Or 5.00 bars